OpenStudy (anonymous):

How do i do this problem? Find the area of a sector with a central angle of 180 degrees and a diameter of 5.6cm. Round to the nearest tenth.

OpenStudy (anonymous):

central angle is180º---> is the area of semicircle \[\frac{A}{2}=\frac{\pi\times r^2}{2}=\frac{\pi(5.6\;\textrm{cm})^2}{2}=49,6\;\textrm{cm}^2\]